# The Effect of Personalization in FedProx: A Fine-grained Analysis on Statistical Accuracy and Communication Efficiency

### v2

@@ A.3 Proof of local statistical Error Bound
$\begin{split}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|_{2}&\leq\frac{1}{2\mu}\left[\frac{\rho}{\sqrt{n}}+\left(\frac{\rho^{2}}{n}+\frac{2\mu\rho^{2}}{nR^{2}}\left(U_{0}+R\right)^{2}\right)^{\frac{1}{2}}\right].\\ &=\frac{1}{\sqrt{n}}\frac{1}{2\mu}\left(\rho+\left(\rho^{2}+\frac{2\mu\rho^{2}}{R^{2}}(U_{0}+R)^{2}\right)^{\frac{1}{2}}\right)\\ &\leq\frac{1}{\sqrt{n}}\frac{\rho+(\rho^{2}+4\mu\rho^{2}(C_{2}+1))^{\frac{1}{2}}}{2\mu}\\ &\leq\frac{1}{\sqrt{n}}\frac{\rho(1+(1+4\mu(C_{2}+1))^{\frac{1}{2}})}{2\mu}.\end{split}$ (61)
Case 2: On the other hand, we will focus on the case that $R$ is relatively small, which means $R\leq\sqrt{\frac{m}{N}}$. From (56), note the definition of $\bm{\delta}_{\star}^{(i)}$ in (20)
$\begin{split}0&\geq-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\left\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}+\delta_{\star}^{(i)}+{{\bm{w}}_{\star}^{(i)}}-{\tilde{\bm{w}}^{(i)}}\right\|^{2}\\ &-\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.\end{split}$ (62)
Note
$2\left\|-\tilde{\Delta}^{(g)}_{\text{stat}}+\tilde{\Delta}^{(i)}_{\text{stat}}+\delta_{\star}^{(i)}\right\|^{2}+2\left\|\tilde{\Delta}^{(g)}_{\text{stat}}-\delta_{\star}^{(i)}\right\|^{2}\geq\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}.$ (63)
Thus
$\left\|-\tilde{\Delta}^{(g)}_{\text{stat}}+\tilde{\Delta}^{(i)}_{\text{stat}}+\delta_{\star}^{(i)}\right\|^{2}\geq\frac{1}{2}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\left\|\tilde{\Delta}^{(g)}_{\text{stat}}-\delta_{\star}^{(i)}\right\|^{2}.$ (64)
Furthermore, note
$\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}+\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}=\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}+\|\bm{\delta}_{\star}^{(i)}\|^{2}\geq\frac{1}{2}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.$ (65)
Plugging (64) and (66) into (62), it yields
@@
$0\geq-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{4}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\lambda}{2}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}-\delta_{\star}^{(i)}\right\|^{2}-\lambda\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}-\lambda\|\bm{\delta}_{\star}^{(i)}\|^{2},$ (66)
which implies
$0\geq-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{4}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-2\lambda\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|^{2}-2\lambda\|\bm{\delta}_{\star}^{(i)}\|^{2}.$ (67)
Apply Assumption 1 and (57) to (67)
$0\geq\left(\mu+\frac{\lambda}{4}\right)\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}-2\lambda U_{0}^{2}-2\lambda R^{2}.$ (68)
Solving (68) as a function of $\|\tilde{\Delta}^{(i)}_{\text{stat}}\|_{2}$, it yields
$\begin{split}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}&\leq\frac{1}{2\mu+\lambda/2}\left\{\frac{\rho}{\sqrt{n_{i}}}+\left[\frac{\rho^{2}}{n_{i}}+4\left(\mu+\frac{\lambda}{4}\right)\lambda\left(U_{0}^{2}+R^{2}\right)\right]^{\frac{1}{2}}\right\}.\end{split}$ (69)
If we assume data size is same among different clients, i.e. $n=n_{i},\forall i\in[m]$ and choose $\lambda=\frac{\rho}{R}\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}=\frac{\rho}{R}\frac{1}{\sqrt{n}}$, (69) can be reformulated as
$\begin{split}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}&\leq\frac{2}{4\mu+\frac{\rho}{\sqrt{n}R}}\left\{\frac{\rho}{\sqrt{n}}+\left[\frac{\rho^{2}}{n}+\left(4\mu+\frac{\rho}{\sqrt{n}R}\right)\frac{\rho}{\sqrt{n}R}\left(U_{0}^{2}+R^{2}\right)\right]^{\frac{1}{2}}\right\}.\end{split}$ (70)
Therefore, we have
$\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}\leq\frac{16\rho^{2}}{\left((4\mu+\frac{\rho}{R\sqrt{n}})\sqrt{n}\right)^{2}}+\frac{8}{4\mu+\frac{\rho}{R\sqrt{n}}}\frac{\rho}{R\sqrt{n}}\left(\frac{C_{1}}{N}+\frac{C_{2}R}{\sqrt{n}}+R^{2}\right).$ (71)
If $\mu<\frac{\rho}{R\sqrt{n}}$
$\frac{\rho}{(4\mu+\frac{\rho}{R\sqrt{n}})\sqrt{n}}\leq\frac{\rho}{\frac{\rho}{R\sqrt{n}}\sqrt{n}}\leq R.$ (72)
If $\mu\geq\frac{\rho}{R\sqrt{n}}$
$\frac{\rho}{(4\mu+\frac{\rho}{R\sqrt{n}})\sqrt{n}}\leq\frac{\rho}{\mu\sqrt{n}}.$ (73)
Combining the discussion in (72) and (73), it yields
$\frac{16\rho^{2}}{\left((4\mu+\frac{\rho}{R\sqrt{n}})\sqrt{n}\right)^{2}}\leq 16\min{\{R^{2},\frac{\rho^{2}}{\mu^{2}n}\}}.$ (74)
Then for the second term on the RHS of (71)
$\begin{split}\frac{8}{4\mu+\frac{\rho}{R\sqrt{n}}}\frac{\rho}{R\sqrt{n}}\left(\frac{C_{1}}{N}+\frac{C_{2}R}{\sqrt{n}}+R^{2}\right)\leq\frac{8C_{1}}{N}+\frac{8C_{2}R}{\sqrt{n}}+\frac{8\rho R}{4\mu\sqrt{n}}.\end{split}$ (75)
Plugging (74) and (75) into (71), we obtain
$\begin{split}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}&\leq 16\min{\{R^{2},\frac{\rho^{2}}{\mu^{2}n}\}}+\frac{8C_{1}}{N}+\frac{8C_{2}R}{\sqrt{n}}+\frac{8\rho R}{4\mu\sqrt{n}}\\ &\leq 16\min{\{R^{2},\frac{\rho^{2}}{\mu^{2}n}\}}+\frac{8C_{1}}{N}+(8C_{2}+\frac{2\rho}{\mu})\frac{R}{\sqrt{n}}\end{split}$ (76)
Combining two-case arguments in (61) and (76), we have:
@@
$\Rightarrow\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\leq\begin{cases}\frac{1}{n}\left(\frac{\rho(1+(1+4\mu(C_{2}+1))^{\frac{1}{2}})}{2\mu}\right)^{2}&R>\sqrt{\frac{m}{N}}\\ 16(R^{2}\wedge\frac{\rho^{2}}{\mu^{2}n})+\frac{8C_{1}}{N}+(8C_{2}+\frac{2\rho}{\mu})\frac{R}{\sqrt{n}}&R\leq\sqrt{\frac{m}{N}}\end{cases}.$ (77)
@@
$\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\leq\frac{C_{3}}{N}+16(R^{2}\wedge\frac{\rho^{2}}{\mu^{2}n})+C_{4}(\frac{R}{\sqrt{n}}\wedge\frac{1}{n}),$ (78)
@@
$C_{3}$ $=8C_{1},$ (79)
$C_{4}$ $=\left(\frac{\rho(1+(1+4\mu(C_{2}+1))^{\frac{1}{2}})}{2\mu}\right)^{2}\wedge(8C_{2}+\frac{2\rho}{\mu}),$ (80)
and $C_{1},C_{2}$ are specified in (47). ∎
@@ Appendix B Optimization Convergence Rate
B.1 Proof of lemmas
B.1.1 Proof of Lemma 2
@@
Proof.
First, from mishchenko2023convergence, we can directly obtain the analytic properties of $F_{i}$. Next, the analytic properties of $h_{i}$ are direct results from Assumption 2 and 4. Finally, to analyze the property of the mapping $w_{\star}^{(i)}$, we just need to use the result from the above claim as we know the definition of ${\bm{w}}_{\star}^{(i)}(\cdot)$ in (8).
B.1.2 Inner Loop Convergence
@@
Suppose Assumption 2 and 4 hold, then if $\eta\leq 1/L_{\ell}$, for all $t\geq 0$ and $\epsilon>0$, the inner loop error bound of all clients can be formulated as:
@@
The proof can be found in Lemma 2 (Ji et al., 2022).
B.1.3 Outer Loop convergence
@@
Suppose Assumption 2 and 4 hold, then if $\gamma\leq L_{g}$, we have for all $t>0$ and $\epsilon_{g}>0$:
@@
$\begin{split}&\|{\bm{w}}_{t+1}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}\\ \leq&(1+\epsilon_{g}\gamma)\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}+\gamma^{2}\|\widetilde{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}+\left(\frac{\gamma}{\epsilon_{g}}\right)\|\widetilde{\nabla}F({\bm{w}}_{t}^{(g)})-\nabla F({\bm{w}}_{t}^{(g)})\|^{2}\\ &-2\gamma(F({\bm{w}}_{t+1}^{(g)})-\widehat{F})-\gamma\mu_{g}\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}\\ &-\gamma^{2}\|\nabla F({\bm{w}}_{t}^{(g)})\|^{2}-\gamma\left(\gamma-L_{g}\gamma^{2}\right)\|\widetilde{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}+\gamma^{2}\|\nabla F({\bm{w}}_{t}^{(g)})-\widetilde{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}\\ \leq&(1-\gamma\mu_{g}+\gamma\epsilon_{g})\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}-\gamma^{2}\|\nabla F({\bm{w}}_{t}^{(g)})\|^{2}+(\gamma^{3}L_{g})\|\widetilde{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}\\ &+\left(\frac{\gamma}{\epsilon_{g}}+\gamma^{2}\right)\|\nabla F({\bm{w}}_{t}^{(g)})-\widetilde{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}.\end{split}$ (95)
B.2 Proof of Theorem 2

### v3

@@ A.3.3 Proof of Local Statistical Error Bound in Theorem 1
$\begin{split}\mathbb{E}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|_{2}&\leq\frac{1}{\mu}\left[\frac{\rho}{\sqrt{n}}+\left(\frac{\rho^{2}}{n}+\frac{2\rho^{2}}{nR^{2}}\left(U_{0}+R\right)^{2}\right)^{\frac{1}{2}}\right].\\ &=\frac{1}{\sqrt{n}}\frac{1}{\mu}\left(\rho+\left(\rho^{2}+\frac{2\rho^{2}}{R^{2}}(U_{0}+R)^{2}\right)^{\frac{1}{2}}\right)\\ &\leq\frac{1}{\sqrt{n}}\frac{\rho+(\rho^{2}+4\rho^{2}(C_{2}+1))^{\frac{1}{2}}}{\mu}\\ &=\frac{1}{\sqrt{n}}\frac{\rho(1+(1+4(C_{2}+1))^{\frac{1}{2}})}{\mu}.\end{split}$ (71)
Case 2: we then consider the case when $R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$.
$\mathbb{E}[\|{\tilde{\bm{w}}^{(i)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}]$
$\leq 3\mathbb{E}[\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|^{2}]+3\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+3\mathbb{E}[\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}]$
$\overset{(a)}{\leq}\frac{12}{(\mu+\lambda)^{2}}\mathbb{E}[\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}})\|^{2}]+3\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+3R^{2}$
@@
$\overset{(b)}{\leq}\frac{12}{(\mu+\lambda)^{2}}\left[4L^{2}\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right]+3\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+3R^{2}$
$\overset{(c)}{\leq}\frac{12}{(\mu+\lambda)^{2}}\left[128L^{2}\left(\frac{\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{L^{2}}{\mu^{2}}R^{2}\right)+4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right]+96\left(\frac{\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{L^{2}}{\mu^{2}}R^{2}\right)+3R^{2}$
$=\frac{12}{(\mu+\lambda)^{2}}\left[\frac{128L^{2}\rho^{2}}{\mu^{2}}\frac{1}{N}+(128L^{4}/\mu^{2}+4L^{2})R^{2}+2\frac{\rho^{2}}{n}\right]+\frac{96\rho^{2}}{\mu^{2}}\frac{1}{N}+(96\frac{L^{2}}{\mu^{2}}+3)R^{2}$ (72)
where step (a) comes from Assumption 1 and Lemma 4, step (b) comes from result (A.3.2) and step (c) is a direct result of applying the error bound of $\mathbb{E}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}$.
If we assume $\lambda$ satisfies the following condition
$\frac{12}{(\mu+\lambda)^{2}}\leq\frac{96\rho^{2}/(\mu^{2}N)+99\kappa^{2}R^{2}}{128\rho^{2}/N+132L^{2}R^{2}+2\rho^{2}/n}$ (73)
we have
$\mathbb{E}[\|{\tilde{\bm{w}}^{(i)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}]\leq\frac{192\rho^{2}}{\mu^{2}}\frac{1}{N}+198\kappa^{2}R^{2}$ (74)
Furthermore, we can get a simplified condition for $\lambda$ as
$\lambda\geq\sqrt{\frac{8(64\rho^{2}/N+66L^{2}R^{2}+\rho^{2}/n)}{32\rho^{2}/(\mu^{2}N)+33\kappa^{2}R^{2}}}-\mu$ (75)
Combining two-case arguments, we have:
@@
$\Rightarrow\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\leq\begin{cases}\frac{1}{n}\left(\frac{\rho(1+(1+4(C_{2}+1))^{\frac{1}{2}})}{\mu}\right)^{2}&R>\sqrt{\frac{m}{N}}\\ \frac{192\rho^{2}}{\mu^{2}}\frac{1}{N}+198\kappa^{2}R^{2}&R\leq\sqrt{\frac{m}{N}}\end{cases}.$ (76)
@@
$\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\leq C_{3}\frac{1}{N}+C_{4}(R^{2}\wedge\frac{1}{n}),$ (77)
@@
$C_{3}$ $=\frac{192\rho^{2}}{\mu^{2}},$ (78)
$C_{4}$ $=\left(\frac{\rho(1+(4C_{2}+5)^{\frac{1}{2}})}{\mu}\right)^{2}\vee 198\kappa^{2},$ (79)
and $C_{2}$ are specified in (57).
Combining the results in A.3.2 and A.3.3, we leave a remark below to further interpret the effect of personalization in obj (1) and discuss how to incorporate the conditions of $\lambda$ when establishing the global and local statistical accuracy.
Remark 2.
To better understand how such an interpolation is achieved through an adaptive personalization degree, we examine the role of $\lambda$ in navigating between the two extreme cases: GlobalTrain and LocalTrain. As the statistical heterogeneity $R\to\infty$, we have $\lambda\to 0$, leading to a high degree of personalization. The statistical error bound becomes $\mathcal{O}(n^{-1})$, matching the rate of LocalTrain; as the statistical heterogeneity $R\to 0$, $\lambda$ increases, leading to a low degree of personalization. The statistical error bound eventually converges to $\mathcal{O}(N^{-1})$, matching the rate of GlobalTrain under a homogeneous setting. As $R\in(0,\infty)$, $\lambda$ transits between the two extremes, making FedProx perform no worse than LocalTrain while benefiting from global training when clients share similarities.
To make the global statistical error bound and local statistical error bound hold simultaneously, we should be aware about the condition of the adaptive strategy of $\lambda$ specified in our proof. It could be summarized as
$\begin{cases}\lambda\geq max(a_{1},a_{2})&R<\sqrt{\frac{m}{N}}\\ \lambda\leq\frac{\rho^{2}}{n\mu R^{2}}&R\geq\sqrt{\frac{m}{N}}\end{cases},$ (81)
where
$a_{1}=\frac{8\kappa}{\mu}\left\{8L^{2}\vee\left(\frac{2L^{2}R^{2}+\rho^{2}/n}{4\rho^{2}/(\mu^{2}N)+4\kappa^{2}R^{2}}\right)\right\}-\mu,$
and
$a_{2}=\sqrt{\frac{8(64\rho^{2}/N+66L^{2}R^{2}+\rho^{2}/n)}{32\rho^{2}/(\mu^{2}N)+33\kappa^{2}R^{2}}}-\mu.$
We can induce a sufficient condition of $\lambda$ for $R<\frac{1}{\sqrt{n}}$ as
$\lambda\geq\text{max}\left\{64\kappa^{2}L,\left(2\kappa\vee 5\right)\mu\frac{2L^{2}R^{2}+\rho^{2}/n}{L^{2}R^{2}+\rho^{2}/N}\right\}-\mu.$ (82)
@@ Appendix B Optimization Convergence Rate
B.1 Algorithm
for $t=1,\dots,T$ do
       The server sends ${\bm{w}}_{t}^{(g)}$ to client $i$, $\forall i\in[m]$
       Set ${\bm{w}}_{t,0}^{(i)}={\bm{w}}_{t-1,K}^{(i)}$;
       for $k=0,\dots,K-1$ do
Algorithm 1 One-stage FedProx
B.2 Proof of lemmas
B.2.1 Proof of Lemma 2
@@
First, from mishchenko2023convergence, we can directly obtain the analytic properties of $F_{i}$. Next, the analytic properties of $h_{i}$ are direct results from Assumption 3 and 2. Finally, to analyze the property of the mapping $w_{\star}^{(i)}$, we just need to use the result from the above claim as we know the definition of ${\bm{w}}_{\star}^{(i)}(\cdot)$ in (9).
B.2.2 Inner Loop Convergence
@@
Suppose Assumption 3 and 2 hold, then if $\eta\leq 1/L_{\ell}$, for all $t\geq 0$ and $\epsilon>0$, the inner loop error bound of all clients can be formulated as:
@@
The proof can be found in Lemma 2 (Ji et al., 2022). ∎
B.2.3 Outer Loop convergence
@@
Suppose Assumption 3 and 2 hold, then if $\gamma\leq L_{g}$, we have for all $t>0$ and $\epsilon_{g}>0$:
@@
$\begin{split}&\|{\bm{w}}_{t+1}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}\\ \leq&(1+\epsilon_{g}\gamma)\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}+\gamma^{2}\|\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}+\left(\frac{\gamma}{\epsilon_{g}}\right)\|\widehat{\nabla}F({\bm{w}}_{t}^{(g)})-\nabla F({\bm{w}}_{t}^{(g)})\|^{2}\\ &-2\gamma(F({\bm{w}}_{t+1}^{(g)})-\widehat{F})-\gamma\mu_{g}\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}\\ &-\gamma^{2}\|\nabla F({\bm{w}}_{t}^{(g)})\|^{2}-\gamma\left(\gamma-L_{g}\gamma^{2}\right)\|\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}+\gamma^{2}\|\nabla F({\bm{w}}_{t}^{(g)})-\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}\\ \leq&(1-\gamma\mu_{g}+\gamma\epsilon_{g})\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}-\gamma^{2}\|\nabla F({\bm{w}}_{t}^{(g)})\|^{2}+(\gamma^{3}L_{g})\|\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}\\ &+\left(\frac{\gamma}{\epsilon_{g}}+\gamma^{2}\right)\|\nabla F({\bm{w}}_{t}^{(g)})-\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}.\end{split}$ (96)
B.3 Proof of Theorem 2